# MLaKE: Multilingual Knowledge Editing Benchmark for Large Language Models

### v1

@@ Abstract
The extensive utilization of large language models (LLMs) underscores the crucial necessity for precise and contemporary knowledge embedded within their intrinsic parameters. Existing research on knowledge editing primarily concentrates on monolingual scenarios, neglecting the complexities presented by multilingual contexts and multi-hop reasoning. To address these challenges, our study introduces MLaKE (Multilingual Language Knowledge Editing), a novel benchmark comprising 4072 multi-hop and 5360 single-hop questions designed to evaluate the adaptability of knowledge editing methods across five languages: English, Chinese, Japanese, French, and German. MLaKE aggregates fact chains from Wikipedia across languages and utilizes LLMs to generate questions in both free-form and multiple-choice. We evaluate the multilingual knowledge editing generalization capabilities of existing methods on MLaKE. Existing knowledge editing methods demonstrate higher success rates in English samples compared to other languages. However, their generalization capabilities are limited in multi-language experiments. Notably, existing knowledge editing methods often show relatively high generalization for languages within the same language family compared to languages from different language families. These results underscore the imperative need for advancements in multilingual knowledge editing and we hope MLaKE can serve as a valuable resource for benchmarking and solution development. [^1] Our benchmark and source code are available at https://github.com/Hi-archers/MLaKE/.
@@ 1 Introduction
With the widespread deployment of large language models, ensuring that the knowledge stored in their intrinsic parameters is correct and up-to-date has became a very important topic (Sinitsin et al., 2020). knowledge editing serves as a promising solution to this challenge, necessitating timely updates to the knowledge embedded within LLMs (Zhu et al., 2020; De Cao et al., 2021; Ding et al., 2024; Meng et al., 2022; Mitchell et al., 2022a; Xu et al., 2024).
@@
Despite considerable efforts devoted to this research field, current studies on knowledge editing typically concentrate on monolingual scenarioi (Li et al., 2023; Zhang et al., 2024; Huang et al., 2023a), where language models are edited and evaluated within the same language (Meng et al., 2022, 2023; Mitchell et al., 2022a). However, the rapid advancements in large language models (LLMs) have facilitated the widespread adoption of multi-lingual settings (Zhao et al., 2023; Wang et al., 2023a). Given this context, the performance of a source-language edited model on other languages remains largely unexplored. Furthermore, as the world evolves, so does the interconnectedness of facts, prompting the question of how effectively an edited model can engage in multi-hop reasoning, wherein the answer to a query relies on a series of interrelated facts rather than a single isolated piece of information (Zhong et al., 2023). These considerations underscore the imperative for these models to adeptly navigate the complexities of multi-hop knowledge across various languages.
@@
To address this challenge, our study introduces MLaKE (Multi-lingual Language Knowledge Editing), a novel benchmark for multilingual multi-hop knowledge editing. MLaKE comprises 5360 single-hop questions and 4072 multi-hop questions designed to test the adaptability of knowledge editing methods across various languages. To ensure the quality and currency of knowledge, we begin by collecting fact chains across languages from Wikipedia. Subsequently, we leverage powerful LLM (e.g., ChatGPT) to generate questions in both free-form and multiple-choice formats using fact chains as input. Consequently, The MLaKE dataset comprises single-hop and multi-hop questions in English, Chinese, Japanese, French, and German, encompassing two distinct formats. This diverse dataset serves as a robust foundation for evaluating the effectiveness of knowledge editing techniques in diverse linguistic environments.
We evaluate the effectiveness of cutting-edge knowledge editing methods on MLaKE to assess their performance in multilingual scenarios. To our surprise, we encounter significant challenges with existing techniques when attempting to edit knowledge in the source language and provide accurate responses in a different language when dealing with single-hop questions, let alone more complicated multi-hop queries. This highlights the substantial impact that language gaps can have on the efficiency of knowledge editing processes. We also conduct in-depth analysis to explore the impact of differences between languages on knowledge editing performance.
Our study not only identifies the limitations of current knowledge editing methods in a multilingual context but also lays the groundwork for future advancements in this area. We hope the MLaKE dataset can serve as a valuable resource for future research, offering a benchmark for developing and testing multilingual knowledge editing solutions. The main contributions of our work are as follows:
@@
  • We evaluate existing knowledge editing methods on MLaKE. Our findings reveal that it remains challenging for multi-lingual LLMs to generalize the edited knowledge to other languages.
  • We observed that existing knowledge editing methods demonstrate superior language generalization within the same language family.
@@ 2 MLaKE: Multi-Lingual Knowledge Editing Benchmark
In our study, we construct the MLaKE (Multi-Lingual Knowledge Editing) benchmark, an comprehensive and challenging dataset that encompasses five languages (English, Chinese, Japanese, French, German), two QA formats (free-form and multiple-choice), and intricate logical structures (single-hop and multi-hop). In this section, we first present the data construction process of MLaKE, followed by a detailed description of the dataset. Lastly, we elaborate on the evaluation settings and metrics utilized.
@@ 2.1 Data Construction of MLaKE
Figure 1 illustrates the construction process of MLaKE, which is primarily composed of three sequential steps: selection and alignment of fact chains, generation of raw data, and construction of QA and MCQA.
@@ 2.1.1 Select chains of facts
We define fact chains as tuples, where a single-hop fact chain is represented as $(s_{1},r_{1},o_{1})$. In this representation, $s_{1}$, $r_{1}$ and $o_{1}$ denote the subject, relationship, and object of the single-hop, respectively. Similarly, a multi-hop fact chain is expressed as $(s_{1},r_{1},o_{1},r_{2},o_{2})$, where $r_{2}$ represents the multi-hop relationship, and $o_{2}$ signifies the second-hop object. Notably, the single-hop object is equivalent to the second-hop subject.
We gather fact chains by crawling data from Wikipedia [^2] We collect data from Wikipedia via the Wikidata API: https://www.wikidata.org/w/api.php. Initially, we manually create a relational dataset consisting of 43 common relations, the same as the approach taken in previous work (Petroni et al., 2019; Meng et al., 2022). Subsequently, we collect single-hop fact chains from Wikidata across five languages, leveraging the relational dataset. During this process, we employ rules to ensure that the single-hop fact chain satisfies specific predefined conditions. For instance, to facilitate batch editing, we enforce restrictions that prevent repetitive modifications of the relationship associated with an entity in the single-hop fact chains. For additional filtering rules applied to fact chains, please refer to the Appendix A.1. To assess the generalizability of the knowledge editing method across languages, we perform alignment on the collected single-hop fact chains. Specifically, we only retain single-hop fact chains that have 5 language forms simultaneously. In addition, we continue to collect knowledge from Wikidata for the objects in the single-hop fact chain to form a multi-hop fact chain. Figure 1 provides a simple example showcasing this process.
Figure 1: Construction of MLaKE. Firstly, we gather a set of common relations and utilize them to extract single-hop and multi-hop fact chains from Wikidata, encompassing five languages. Then, we combine ChatGPT and manual collaboration to generate edited objects for them, and align the single-hop fact chains. Finally, we utilize the organized raw data to create two types of samples: QA format and MCQA format.
@@ 2.1.2 Generation of raw data
2.1.3 Construction of QA and MCQA
To cater to various models, we generate both free-form and multiple-choice questions for all samples in the raw data. It is important to highlight that the questions in the corresponding QA and MCQA samples are identical, differing only in their format. The MCQA format includes three additional options. Refer to Appendix  A for comprehensive instructions.
During the aforementioned process, in addition to utilizing ChatGPT, we conduct necessary manual verification and implemented fixed rules to ensure data quality, see the Appendix A.1.

### v2

@@ Abstract
The extensive utilization of large language models (LLMs) underscores the crucial necessity for precise and contemporary knowledge embedded within their intrinsic parameters. Existing research on knowledge editing primarily concentrates on monolingual scenarios, neglecting the complexities presented by multilingual contexts and multi-hop reasoning. To address these challenges, our study introduces MLaKE (Multilingual Language Knowledge Editing), a novel benchmark comprising 4072 multi-hop and 5360 single-hop questions designed to evaluate the adaptability of knowledge editing methods across five languages: English, Chinese, Japanese, French, and German. MLaKE aggregates fact chains from Wikipedia across languages and utilizes LLMs to generate questions and answer. We assessed the effectiveness of current multilingual knowledge editing methods using the MLaKE dataset. Our results show that due to considerable inconsistencies in both multilingual performance and encoding efficiency, these methods struggle to generalize effectively across languages. The accuracy of these methods when editing English is notably higher than for other languages. The experimental results further demonstrate that models encode knowledge and generation capabilities for different languages using distinct parameters, leading to poor cross-lingual transfer performance in current methods. Transfer performance is notably better within the same language family compared to across different families. These findings emphasize the urgent need to improve multilingual knowledge editing methods. [^1] Our benchmark and source code are available at https://anonymous.4open.science/r/MLaKE_an-4D96/.
MLaKE: Multilingual Knowledge Editing Benchmark for
Large Language Models
[…]
@@ 1 Introduction
With the widespread deployment of large language models, ensuring that the knowledge stored in their intrinsic parameters is correct and up-to-date has become a very important topic (Sinitsin et al., 2020; Chen and Shu, 2023, 2024). knowledge editing serves as a promising solution to this challenge, necessitating timely updates to the knowledge embedded within LLMs (Zhu et al., 2020; De Cao et al., 2021; Ding et al., 2024; Meng et al., 2022; Mitchell et al., 2022a; Xu et al., 2024; Chen et al., 2024).
@@
Despite considerable efforts devoted to this research field, current studies on knowledge editing typically concentrate on monolingual scenarioi (Li et al., 2023; Zhang et al., 2024; Huang et al., 2023a), where language models are edited and evaluated within the same language (Meng et al., 2022, 2023; Mitchell et al., 2022a). However, the rapid advancements in large language models (LLMs) have facilitated the widespread adoption of multi-lingual settings (Zhao et al., 2023; Wang et al., 2023a). Given this context, the performance of a source-language edited model on other languages remains largely unexplored.
@@
To address this challenge, our study introduces MLaKE (Multi-lingual Language Knowledge Editing), a novel benchmark for multilingual multi-hop knowledge editing. MLaKE comprises 5360 single-hop questions and 4072 multi-hop questions designed to test the adaptability of knowledge editing methods across various languages. To ensure the quality and currency of knowledge, we begin by collecting fact chains across languages from Wikipedia. Subsequently, we leverage powerful LLM (e.g., ChatGPT) to generate questions in both free-form and multiple-choice formats using fact chains as input. Consequently, The MLaKE dataset comprises single-hop and multi-hop questions in English, Chinese, Japanese, French, and German. This diverse dataset serves as a robust foundation for evaluating the effectiveness of knowledge editing techniques in diverse linguistic environments.
We assessed the effectiveness of various knowledge editing methods on MLaKE, with a focus on their performance in multilingual contexts. The results show that current mainstream knowledge editing methods demonstrate weak generalization in multilingual editing and poor cross-lingual transfer performance. Specifically, these methods not only struggle to edit knowledge in one language while transferring across others, but also exhibits weak performance in the foundational task of editing knowledge within a single language. Moreover, the aforementioned challenges become even more pronounced in multi-hop reasoning scenarios. Our findings underscore the significant impact of language differences on the performance of knowledge editing. To better understand this challenge, we conduct a series of experiments to analyze the effects of linguistic and structural differences on knowledge editing performance.
The main contributions of our work are as:
@@
  • We demonstrate that existing knowledge editing methods, when applied to LLMs, suffer from significant shortcomings in multilingual generalization and cross-lingual transfer performance.
  • Our analysis shows that weak multilingual generalization is primarily due to the models’ insufficient multilingual performance and encoding inefficiency. The limited cross-lingual transferability is largely caused by LLMs using distinct parameter sets to encode knowledge for different languages.
@@ 2 MLaKE: Multi-Lingual Knowledge Editing Benchmark
Figure 1: Construction of MLaKE. Firstly, we gather a set of common relations and utilize them to extract single-hop and multi-hop fact chains from Wikidata, encompassing five languages. Then, we combine ChatGPT and manual collaboration to generate edited objects for them, and align the single-hop fact chains. Finally, we utilize the organized raw data to create QA data.
In our study, we construct the MLaKE (Multi-Lingual Knowledge Editing) benchmark, an comprehensive and challenging dataset that encompasses five languages (English, Chinese, Japanese, French, German) and intricate logical structures (single-hop and multi-hop). In this section, we first present the data construction process of MLaKE, followed by a detailed description of the dataset. Lastly, we elaborate on the evaluation settings and metrics utilized.
@@ 2.1 Data Construction of MLaKE
Figure 1 illustrates the construction process of MLaKE, which is primarily composed of three sequential steps: selection and alignment of fact chains, generation of raw data, and construction of question and answer.
@@ 2.1.1 Select chains of facts
We define fact chains as tuples, where a single-hop fact chain is represented as $(s_{1},r_{1},o_{1})$. In this representation, $s_{1}$, $r_{1}$ and $o_{1}$ denote the subject, relationship, and object of the single-hop, respectively. Similarly, a multi-hop fact chain is expressed as $(s_{1},r_{1},o_{1},r_{2},o_{2})$, where $r_{2}$ represents the multi-hop relationship, and $o_{2}$ signifies the multi-hop object. Notably, the single-hop object is equivalent to the multi-hop subject.
Inspired by the work of Zhong et al. (2023), We gather fact chains by crawling data from Wikipedia [^2] We collect data from Wikipedia via the Wikidata API: https://www.wikidata.org/w/api.php. Initially, we manually create a relational dataset consisting of 43 common relations, the same as the approach taken in previous work (Petroni et al., 2019; Meng et al., 2022). Subsequently, we collect single-hop fact chains from Wikidata across five languages, leveraging the relational dataset. During this process, we employ rules to ensure that the single-hop fact chain satisfies specific predefined conditions. For instance, to facilitate batch editing, we enforce restrictions that prevent repetitive modifications of the relationship associated with an entity in the single-hop fact chains. For additional filtering rules applied to fact chains, please refer to the Appendix C.1. To assess the generalizability of the knowledge editing method across languages, we perform alignment on the collected single-hop fact chains. Specifically, we retain only single-hop fact chains that were simultaneously available in all five languages. Given that Wikipedia is written by local communities worldwide, this approach allow us to gather authentic localized expressions. In addition, we continue to collect knowledge from Wikidata for the objects in the single-hop fact chain to form a multi-hop fact chain. Figure 1 provides a simple example showcasing this process.
@@ 2.1.2 Generation of raw data
2.1.3 Construction of question and answer
We employ ChatGPT to generate coherent and fluent multilingual questions, along with corresponding answers, from the collected triplet data in various languages. Instead of using the commonly applied template-based approach from previous research to convert triplets into questions, we opt for a different method due to the complexity of word inflections in French and German sentence structures. These complexities make natural conversion difficult when relying solely on template filling. To minimize potential errors in this transformation process, five experts in the relevant languages were invited to review the model-generated texts, following the criteria detailed in Appendix C.1.
@@ 2.2 Description of the MLaKE
Dataset statistics
Table 5 summarizes the statistics of the MLaKE dataset. The MLaKE dataset consists of more than 13K samples. We align all single-hop problems and employ them as a means to investigate the generalizability of existing knowledge editing methods across different languages following editing in a single language. Multi-hop problems are not aligned across different languages, and we employ them to assess the generalizability and portability of existing editing methods.
Dataset analysis
Figure 8 briefly analyzes the characteristics of the MLaKE dataset. Figure 8(a) depicts all first relations and their corresponding top 3 second relations, demonstrating the diversity of relations in MLaKE. The majority of questions pertain to nationality, names of individuals and locations, and typically adhere to the following structure: "From which nation does Gwendoline Christie hold citizenship?" (single-hop question) or "In which country is the institution where Carl Sagan was employed located?" (multi-hop question). In Figure 8(b), we further examine the relation PIDs that account for more than 1%, mainly including "country of citizenship" (p27), "country" (P17), "continent" (P30), etc. For the corresponding table of relationship PID and relationship label, please refer to the Appendix C. Figure 8(c) depicts the distribution of entities that have an occurrence rate exceeding 0.5%. The prominent entities in this distribution include ’United Kingdom’ (Q145), ’Canada’ (Q36), and ’United States of America’ (Q30). Figure 8(d) illustrates the distribution of question lengths. The majority of questions fall within the 10-20 word range, which allows for precise expression of the subject and relationship without the inclusion of extraneous information. To accommodate various answer preferences of Large Language Models (LLMs), we strive to generate multiple aliases for each answer in MLaKE. Figure 8(e) demonstrates that the majority of answers possess 2-13 aliases, while there are even several answers with more than 20 aliases. The analysis data presented in Figure 8 only includes English samples, both single-hop and multi-hop.
2.3 Evaluation Metrics
Diverging from other benchmarks, MLaKE primarily focuses on assessing the generalizability and portability of knowledge editing methods across multilingual scenarios. For different models, we use corresponding question to guide them to generate answers. We evaluate the accuracy of the question-answering task by determining whether the model-generated sentences contained the correct answer or its aliases.
3 Experiments
This section first explains the experimental settings, then analyzes the generalization ability of multilingual knowledge editing and the transfer ability of cross-language knowledge editing, and finally explores the potential reasons that affect the performance of knowledge editing.
3.1 Experimental Setup
Language Models
We use the following two language models in our experiments: 1) Vicuna-7B-v1.5 is fine-tuned from Llama-2 using supervised instruction fine-tuning with training data sourced from ShareGPT.com. 2) Qwen1.5-7B-Chat is a transformer-based aligned chat model pre-trained on extensive data, developed by Alibaba Cloud.
Knowledge Editing Methods
Evaluation Dimensions
  • Multilingual generalization refers to the accuracy of knowledge editing across different languages, especially the editing accuracy in other languages after using a single language for editing.
  • Cross-lingual transferability refers to the ability of the editing model to apply the knowledge edited in one language to another language, i.e., the accuracy of editing in other languages after performing editing in one language.
Methods Single-hop QA (%) - Vicuna Multi-hop QA (%) - Vicuna
EN DE FR JA ZH EN DE FR JA ZH
MEND 0.19 0.09 0.09 0.00 0.00 0.00 0.14 0.86 0.04 0.25
ROME 14.93 1.59 4.38 0.09 0.28 2.44 0.00 0.57 0.11 0.25
MEMIT 60.73 23.79 44.22 4.94 3.73 20.70 7.11 14.12 3.19 3.30
StableKE 88.43 37.31 37.31 32.09 28.73 26.66 13.68 5.99 11.11 11.66
Methods Single-hop QA (%) - Qwen Multi-hop QA (%) - Qwen
EN DE FR JA ZH EN DE FR JA ZH
ROME 44.59 21.92 20.24 0.28 33.68 15.72 4.51 5.85 0.11 16.48
MEMIT 75.47 59.51 51.12 20.24 43.38 41.31 19.56 18.12 9.46 27.50
StableKE 92.82 45.34 44.96 37.31 67.44 55.37 22.02 10.56 17.05 31.94
Table 1: Single-hop and Multi-hop QA performance comparison between Vicuna and Qwen models using various knowledge editing methods across different languages.
3.2 Generalization of Multilingual Knowledge Editing
To evaluate the multilingual generalization performance of commonly used knowledge editing methods. We select four widely used knowledge editing methods and evaluated their robustness across five different languages using the Vicuna and Qwen models. Each method are both edited and tested within the same language context. Table 1 presents the experimental results, and the main conclusion are as follows.
Conclusion 1: All evaluated knowledge editing methods exhibit limited multilingual generalization. As can be seen in Table 1, all methods demonstrate significantly higher accuracy in English knowledge editing compared to other languages, regardless of the base model used. This may be due to the fact that the quality and scale of English are the highest among the training corpora used by existing models.
Conclusion 2: The performance of the knowledge editing method is related to the performance of the base model. Except for StableKE, the other three methods perform poorly in editing Chinese and Japanese knowledge using the Vicuna model. In contrast, these methods achieve significantly higher accuracy with the Qwen model. The Figure 2 shows that the better the base model performs in a particular language, the more effective the knowledge editing method becomes.
Figure 2: The relationship between the original performance and the edited performance of the model. With the same knowledge editing method, the better the model’s original performance, the more effective the knowledge editing becomes.
Additionally, the table shows that the accuracy of multi-hop knowledge editing is substantially lower than that of single-hop editing. This finding aligns with previous studies focused on English, and our study extends these conclusions to a multilingual scenario.
Figure 3: Performance of MEMIT and StableKE on different source and edit languages on the vicuna1.5 and Qwen1.5 model.
Figure 4: Relationship between the average number of tokens needed to encode each subject and the success rate of editing.
3.3 Transferability of Cross-Language Knowledge Editing
To evaluate the transferability performance of existing knowledge editing methods in cross-language scenarios, we conduct knowledge editing in one language and assessed the editing accuracy in different languages. Our analysis focus on two representative methods: MEMIT and StableKE. Results from additional methods are presented in the appendix, and the main conclusion are as follows.
Conclusion 3: All knowledge editing methods have limited cross-language transferability, and the accuracy of the model in answering questions drops significantly when using different languages for reasoning. As illustrated in Figures 3, the editing performance along the diagonal is notably superior, indicating that the cross-language generalization ability of MEMIT and StableKE is limited. Notably, when performing knowledge editing on English on the Vicuna model, the accuracy in other languages was significantly lower than in English. A similar pattern was observed in the Qwen model during both Chinese and English editing.
Conclusion 4: The greater the similarity between the editing language and the reasoning language, the stronger the cross-language generalization ability. According to linguistic classification, English, German, and French all belong to the Indo-European language family Joseph (2005). Although Chinese and Japanese do not belong to the same language family, Japanese has been significantly influenced by Chinese Handel (2008); Jin-lian (2004). The figure reflects a trend: in the top-left corner, the cross-linguistic generalization among the three Indo-European languages is notably stronger, whereas generalization from these languages to Chinese and Japanese is significantly weaker than within the Indo-European family. While there is some connection between Chinese and Japanese, it is relatively weak, resulting in stronger cross-linguistic generalization between these two languages compared to their generalization to other languages, yet weaker than the generalization observed within the Indo-European language family.
Method Accuracy
Use Subject No Subject
ROME 14.93 6.16
MEMIT 60.73 4.91
Table 2: Accuracy of MEMIT and ROME with and without using subject.
4 Causes of Multilingual Editing Challenges
Figure 5: Proportion of responses that remain identical to the original, unedited outputs after applying MEMIT and StableKE with different source and edit languages on the Vicuna1.5 and Qwen1.5 model.
We identify two causes contributing to the suboptimal performance of multilingual knowledge editing: significant disparities in the multilingual capabilities of models and the structural independence of multilingual representations. The inconsistency in multilingual performance refers to the model’s varying proficiency and encoding efficiency across different languages, which undermines the generalization of knowledge editing in multilingual contexts. Structural independence in multilingual models suggests that knowledge and abilities in different languages are encoded by distinct parameters. This separation limits the effectiveness of knowledge editing in achieving cross-linguistic transfer.
4.1 Potential Factors Limiting Multilingual Editing Generalization
Besides the model’s multilingual performance, another crucial factor influencing the generalization ability of multilingual knowledge editing methods is its text encoding performance. To examine the relationship between text encoding performance and the generalization of multilingual knowledge editing, we analyze the ROME and MEMIT methods. Both methods follow the locate-then-edit approach, with their core process focused on identifying the subject’s final token to enhance editing accuracy. However, as shown in Table 2, when the subject’s final token is replaced with the sentence’s last token, disrupting this process, the knowledge editing accuracy of both MEMIT and ROME declines significantly. Additionally, the Vicuna and Qwen models exhibit substantial differences in encoding efficiency across languages. For instance, as demonstrated in Table 3, the Vicuna model requires an average of 11.77 tokens to represent a Chinese subject, while only 4.04 tokens are needed for an English subject. As a result, the information density in the final token of a Chinese subject in Vicuna is lower than that of an English subject, negatively impacting target localization accuracy and reducing editing precision. As illustrated in Figure 4, as a model’s encoding efficiency decreases (i.e., more tokens are required to encode a target), the success rate of knowledge editing correspondingly diminishes. In addition to the more apparent factor of the model’s multilingual performance influencing generalization, we further demonstrate through experiments that the model’s multilingual text encoding efficiency is another potential factor limiting multilingual editing generalization.
4.2 Examining Factors Affecting Cross-lingual Knowledge Transfer
Conclusion 5: Knowledge editing has a greater impact on the knowledge of languages that are more closely related to each other. To thoroughly investigate the factors influencing cross-lingual transferability in multilingual knowledge editing, we focus on evaluating how edits to knowledge affect content generated in languages that were not directly edited. we further analyze how MEMIT and StableKE impacted the models’ ability to generate multilingual text. As shown in Figures 5, our results further illustrate the findings presented in Section 3.3. Knowledge editing exerts a stronger influence on the knowledge of languages closely related to the target language, while exerting a weaker influence on the knowledge of more distantly related languages. For example, when knowledge editing is performed in Chinese or Japanese, the accuracy of the original responses in three Indo-European languages decreases only slightly, indicating minimal disruption. These results suggest that existing knowledge editing methods have a smaller effect on knowledge associated with languages that are less closely related. For instance, when knowledge editing is conducted in Chinese or Japanese, the accuracy of responses in three Indo-European languages declines only marginally, indicating minimal interference. Similarly, when Indo-European languages are used for knowledge editing, the accuracy of original responses in Chinese and Japanese shows a comparable marginal decline, reflecting limited impact. These findings suggest that current knowledge editing methods have a reduced impact on knowledge related to languages that are less closely connected.
Model EN DE FR JA ZH
Vicuna 4.04 4.34 4.51 9.72 11.77
Qwen 3.74 4.14 4.29 6.27 5.17
Table 3: Average number of tokens needed to encode a subject word by Vicuna and Qwen.
Figure 6: Fluency Performance of MEMIT and StableKE on different source and edit languages on the Vicuna1.5 and Qwen1.5 model.
Figure 7: Overlap of the Top 100 and 200 Most Impacted Neurons Across Different Languages When Editing Vicuna and Qwen Models Using MEMIT and StableKE.
Conclusion 6: Knowledge editing has a greater impact on the generation ability of languages that are more closely related to each other. To comprehensively evaluate the impact of multilingual knowledge editing on non-edited languages, we assess its impact on the generative ability of non-edited languages. Without considering the accuracy of the text, we used ChatGPT-4 to evaluate the fluency of the content. As presented in Figure 6, knowledge editing minimally affects the generative performance of distantly related languages. Notably, when MEMIT is applied to edit Vicuna’s knowledge in Chinese, the model’s ability to generate coherent Chinese text deteriorated significantly, resulting in repetitive, incoherent, and meaningless output. In contrast, the impact on English text generation was less pronounced. Specifically, the fluency score for Chinese dropped sharply from 3.84 to 2.29, while the English fluency score declined only slightly from 4.03 to 3.81 following knowledge editing. This trend is observed consistently across multiple language families. We hypothesize that the model encodes knowledge and generative abilities for different languages in distinct parameters, with greater divergence in encoding for more distantly related languages.
To test this hypothesis, we analyze the differences in the multilayer perceptrons (MLPs) that are significantly affected at the same layer during knowledge editing across different languages. Specifically, we identify the neurons that exhibited the largest changes when editing the Vicuna and Qwen models using the MEMIT and StaleKE methods. We then compare the extent of overlap in these neurons across various languages. As illustrated in Figure 7, the MEMIT method, which focuses on layers 4-8 for editing, operates within a narrower range but exerts a more pronounced effect on the corresponding parameters. We identified the top 100 neurons most impacted by MEMIT and analyzed their cross-linguistic overlap. Additionally, We analyze the overlap between the top 200 neurons primarily influenced by the StaleKE approach, which modifies all parameters of the model. Although StaleKE affects a broader range of parameters, its influence on individual neurons is comparatively smaller. The results indicate that the cross-linguistic overlap of the most impacted neurons is generally low, with all overlaps remaining below 20%. However, closely related languages demonstrate a higher degree of neuron overlap. For example, languages within the Indo-European family exhibit greater overlap compared to those from different language families. Given that knowledge and capabilities in different languages are encoded in distinct parameters, current knowledge-editing methods face limitations in their ability to generalize edits across languages when applied to a single language.
5 Conclusion
In this paper, we explore the effectiveness of current knowledge editing methods in multilingual settings. To this end, we create the MLaKE dataset by extracting knowledge tuples from Wikipedia and generating single-hop and multi-hop questions using ChatGPT. Leveraging MLaKE, we conducte experiments employing various methods and multilingual LLMs to investigate the transferability of knowledge editing from English to other languages and vice versa. Our research findings reveal that: (1) Current knowledge-editing methods have limited generalization performance in multilingual settings. Beyond the inconsistent model performance across different languages, experiments suggest that insufficient multilingual encoding may contribute to their weak generalization. (2) The cross-lingual transferability of knowledge-editing methods is similarly constrained. Even within the same language family, transferability remains limited and is especially weak across distinct language families. A key finding from the experiments is that LLMs encode knowledge in different languages using distinct parameter sets, which reduces cross-lingual transfer performance.
Limitations
In this paper, we primarily explore knowledge-editing methods, focusing on parameter adjustment. These methods specifically target knowledge encoded in model parameters, and through a comprehensive analysis of their impact on the parameters, we highlight existing limitations in multilingual generalization and cross-lingual knowledge transfer. Our findings provide insights for advancing the development of more robust multilingual knowledge-editing methods. Given time and resource constraints, we do not extensively explore or compare knowledge-editing methods unrelated to parameter changes, like in-context editing.
Ethics Statement
This study thoroughly investigates multilingual knowledge editing in large language models and introduces the Multilingual Knowledge Editing Dataset-MLaKE. Using this dataset, we extensively demonstrate and analyze the performance of existing knowledge editing methods across different languages. The MLaKE dataset reveals notable disparities in current multilingual knowledge editing methods, specifically that the success rate of knowledge editing in English is significantly higher than in other languages. Consequently, MLaKE facilitates the advancement of more equitable and generalizable multilingual knowledge editing methods.
Appendix A Evaluating QA Performance of Five Widely Used LLMs on MLaKE
To evaluate the effectiveness of LLMs in various languages and their content generation capabilities, we assessed five instruction-tuned LLMs: LLaMa-2-7B-chat, Vicuna-7B-v1.5, Qwen1.5-7B-Chat, Gemma-7B-IT, and Mistral-7B-Instruct-v0.2 s Touvron et al. (2023); Team et al. (2024); Bai et al. (2023); Jiang et al. (2023). Using the unedited MLaKE data, we assessed two capabilities of these five models. The results for single-hop and multi-hop QA are presented in Table  4. Our key findings include: 1. The same model can display notable differences in QA performance when tested in different languages. English consistently performs better in both single-hop and multi-hop QA, indicating that LLMs have a deeper grasp of English. 2. Qwen1.5-7B-Chat demonstrates a more consistent performance across various languages than other LLMs, achieving especially strong results in Chinese and Japanese. We attribute this advantage to its pre-training corpus, which includes a diverse range of languages Bai et al. (2023). 3. Conversely, Vicuna-7B-v1.5 shows markedly improved generative capabilities in Chinese and Japanese. This enhancement is mainly attributed to the inclusion of Chinese and Japanese data in the ShareGPT instruction-tuning dataset, which strengthens its ability to generate responses in these languages. 4. Gemma-7B-IT shows significant differences in accuracy in free-form QA compared to other models. Notably, it falls behind LLaMa-2-7B-Chat by 27.8% in single-hop Free form QA and 15.11% in multi-hop Free form QA. This gap is mainly attributed to the RLHF optimization in the Gemma model, which often leads to its refusal to generate responses.
{CJK*}
UTF8gbsn
Models Single-hop QA (%) Multi-hop QA (%)
EN DE FR JA ZH EN DE FR JA ZH
LLaMa-2-7B-Chat 78.17 59.05 63.62 0.0 0.0 49.94 25.99 29.92 0.0 0.0
Vicuna-7B-v1.5 76.21 58.40 61.29 46.83 36.94 47.60 28.50 26.98 26.50 21.32
Qwen1.5-7B-Chat 78.92 54.85 58.21 37.87 57.00 60.81 30.46 30.56 30.51 53.55
Gemma-7B-IT 50.37 42.07 42.35 43.00 34.33 34.83 12.94 22.51 24.26 22.11
Mistral-7B-Ins.-v0.2 83.21 60.73 65.95 26.77 28.17 62.12 40.39 37.21 18.85 29.21
Table 4: Single-hop and Multi-hop Free-form QA performance of five LLMs on MLaKE.
Appendix B Implementation Details
All experiments are conducted on a single NVIDIA A800 GPU (80GB). We re-implement MEND, ROME and MEMIT using EasyEdit [^3] https://github.com/zjunlp/EasyEdit (Wang et al., 2023c) with default settings. Additionally, we reproduce the results of StableKE by utilizing their official repository [^4] https://github.com/Hi-archers/StableKE. In the case of MEND, MEMIT, and StableKE, a batch size of 100 was employed for actual editing. As for ROME, which does not support batch editing, a batch size of 1 was used, and 100 samples were iteratively edited before testing.
Appendix C Data Details